determine all possible values of k so that the quadratic below has one zero.
x^2 + (k-5)x + (4-5k) = 0
the asnwers are k = -1 and k = -9. thank you!

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

This needs completing the square, since that would give one zero with multiplicity 2. This is done by taking the coefficient of the x term, dividing it by 2 and squaring it. That will give the coefficient of the constant.
Here, those values are:
[(1/2)(k-5)]^2=(4-5k)
(1/4)(k^2-10k+25)=(4-5k)
k^2-10k+25=16-20k, multiplying by 4
k^2+10k+9=0
(k+9)(k+1)=0
k=-9 and -1

Answer by ikleyn(52780)

.
A quadratic equation   ax^2 + bx + c = 0


    - has two real zeroes if and only if the discriminant d = b^2 - 4ac is POSITIVE;

    - has one real zero   if and only if the discriminant d = b^2 - 4ac is ZERO;

    - has NO real zeroes if and only if the discriminant d = b^2 - 4ac is NEGATIVE.


So, for your equation, in order for the condition was satisfied, the discriminant 

(k-5)^2 - 4*(4-5k)  must be equal to zero:


(k-5)^2 - 4*(4-5k) = 0.


It is your equation to find the values of "k".


Simplify and solve it by any method you know.

--------------
If you still have questions, you can post them to me through the "Thank you" window/form.

In his case refer to the problem ID number 1106233.

I will help you free of charge.